OpenStudy (anonymous):

Solve for a: `-2a + 10 = -12` A. 1 B. 20 C. 15 D. 11

OpenStudy (akashdeepdeb):

-2a + 10 = -12 2a = 22 a = ? What do you think? :)

OpenStudy (anonymous):

idk :(

OpenStudy (anonymous):

D

OpenStudy (akashdeepdeb):

Yes it is D 11 But did you get why?

OpenStudy (zzr0ck3r):

@jaily18 read the rules don't just give the answers, we are not here to give the answers

OpenStudy (anonymous):

-2A+10=-12 -10 -10 -2A=-22 A=11

OpenStudy (zzr0ck3r):

also you should try and lead them to the question. not just give them the answer and the steps...

OpenStudy (anonymous):

you have to subtract -10 to both sides so that you can get -2A by itself... so -10+(-12)=-22 then write it as -2A=-22, divide -2 to both sides to get A by itself. A=11 because -22/-2 = 11